# Distribution-consistency Structural Causal Models

### 30 Jan 2024 v2

@@ A DiscoSCM is a tuple $\langle U,\mathbf{E},\mathbf{V},\mathcal{F}\rangle$, where
$\mathcal{F}_{\mathbf{x}}:=\{f_{i}:V_{i}\notin\mathbf{X}\}\cup\{\mathbf{X}\leftarrow x\},$
@@ 5.1 Main Theoretical Results
is referred to as Layer 1 valuation. Here, $\mathbf{y}$ and $\mathbf{z}$ represent the observed outcomes, $\mathbf{x}$ and $\mathbf{w}$ the observed treatments, $\mathbf{e}$ the noise instantiation, $u$ the individual, and we denote $\mathbf{y}_{\mathbf{x}}$ and $\mathbf{z}_{\mathbf{w}}$ as the realization of their corresponding potential outcomes, $\mathbf{e}_{\mathbf{x}}$, $\mathbf{e}_{\mathbf{w}}$ as the instantiation of their corresponding counterfactual noises.
@@
Here, $e$ represents the observed trace, also referred to as evidence, exemplified by $X=\tilde{x},Y=\tilde{y}$, concisely denoted as $e=[\tilde{x},\tilde{y}]$. In contrast, to correspondingly conduct Layer 3 valuation $P(Y^{d}(x)=y|e)$ in DiscoSCMs, we have a theorem as follows.
@@
$\stackrel{{\scriptstyle\triangle}}{{=}}\sum_{u}P(y_{x};u)P(u|e)$
The first equality follows the total probability formula, and the second equality is due to Lemma 3 that is derived from Assumption 3. This completes the proof. ∎
@@ 5.2 Illustration with Examples
It is noteworthy that for a SCM describing the same system, Layer 1 and 2 valuations in the SCM are the same as those in the DiscoSCM, as stated in Theorem 1. However, the two Layer 3 valuations within the DiscoSCM, $P(Y^{d}=-1|T=-1)$ and $P(Y^{d}(t=-1)=-1|T=-1,Y=-1)$, exhibit discrepancies when compared to their counterparts in the SCM. Specifically, we observe the following differences:
@@ 5.3 Further Exploration of the Unit Selection Variable
The above two theorems respectively correspond to Theorems 7.1&7.18 in Neal (2020) related to the propensity score in the PO framework. The key difference is that in our theorems, $U$ replaces the pre-treatment features that satisfy the unconfoundedness assumption(Imbens and Rubin, 2015). This suggests that in our DiscoSCM framework, $U$ plays a role analogous to that of pre-treatment features.

### 23 Mar 2024 v3

@@ A DiscoSCM is a tuple $\langle U,\mathbf{E},\mathbf{V},\mathcal{F}\rangle$, where
$\mathcal{F}_{\mathbf{x}}:=\{f_{i}:V_{i}\notin\mathbf{X}\}\cup\{\mathbf{X}\leftarrow\mathbf{x}\},$
@@ 5.1 Main Theoretical Results
is referred to as Layer 1 valuation. Here, $\mathbf{y}$ and $\mathbf{z}$ represent the observed outcomes, $\mathbf{x}$ and $\mathbf{w}$ the observed treatments, $\mathbf{e}$ the noise instantiation, $u$ the individual, and we denote $\mathbf{y}_{\mathbf{x}}$ and $\mathbf{z}_{\mathbf{w}}$ as the realization of their corresponding counterfactual outcomes, $\mathbf{e}_{\mathbf{x}}$, $\mathbf{e}_{\mathbf{w}}$ as the instantiation of their corresponding counterfactual noises.
@@
Here, $e$ represents the observed trace, also referred to as evidence, exemplified by $X=\tilde{x},Y=\tilde{y}$, concisely denoted as $e=[\tilde{x},\tilde{y}]$. [^9] Please distinguish between the “evidence” ($e$) and the instantiation of “exogenous noise ” ($\mathbf{e}$). In contrast, to correspondingly conduct Layer 3 valuation $P(Y^{d}(x)=y|e)$ in DiscoSCMs, we have a theorem as follows.
@@
$\stackrel{{\scriptstyle\triangle}}{{=}}\sum_{u}P(y|x;u)P(u|e)$
The first equality follows the total probability formula, and the second equality is due to Theorem 3 that is derived from Assumption 3. This completes the proof. ∎
@@ 5.2 Illustration with Examples
It is noteworthy that for an SCM describing the same system, Layer 1 and 2 valuations in the SCM are the same as those in the DiscoSCM, as stated in Theorem 1. However, the two Layer 3 valuations within the DiscoSCM, $P(Y^{d}=-1|T=-1)$ and $P(Y^{d}(t=-1)=-1|T=-1,Y=-1)$, exhibit discrepancies when compared to their counterparts in the SCM. Specifically, we observe the following differences:
@@ 5.3 Further Exploration of the Unit Selection Variable
The above two theorems respectively correspond to Theorems 7.1&7.18 in Neal (2020) related to the propensity score in the PO framework. The key difference is that in our theorems, $U$ replaces the pre-treatment features that satisfy the unconfoundedness assumption(Imbens and Rubin, 2015). This suggests that in our DiscoSCM framework, $U$ plays a role analogous to that of pre-treatment features. Moreover, if we assume the conditional independence between $X$ and $(T,Y)$ given $U$ in the underlying causal graph, the following theorems hold.
Theorem 7.
For a DiscoSCM with discrete treatment $T$, feature $X$ and outcome $Y$:
$E[Y(t^{\prime})|X=x^{\prime}]=E[\frac{Y\mathbf{1}_{T=t^{\prime},X=x^{\prime}}}{P(t^{\prime}|U)P(x^{\prime})}]$ (29)
where $t^{\prime}$ is any given treatment value and $P(t^{\prime}|u)$ is positive for any individual $U=u$.
Proof.
On one hand, we have:
$E[Y(t^{\prime})|X=x^{\prime}]$ $=E[E[Y(t^{\prime})|X=x^{\prime},U]|X=x^{\prime}]$
$=\sum_{u}\sum_{y}yP(y_{t^{\prime}}|x^{\prime};u)P(u|x^{\prime})$
$=\sum_{u}\sum_{y}yP(y|{t^{\prime}};u)P(u|x^{\prime})$
$=\sum_{u}\sum_{y}y\frac{P(y|t^{\prime},u)P(t^{\prime}|u)P(u)}{P(t^{\prime}|u)}\cdot\frac{P(x^{\prime}|u,t^{\prime},y)}{P(x^{\prime}|u)}\cdot\frac{P(x^{\prime}|u)}{P(x^{\prime})}$
$=\sum_{u}\sum_{y}y\frac{P(u,t^{\prime},y)}{P(t^{\prime}|u)}\cdot\frac{P(x^{\prime}|u,t^{\prime},y)}{P(x^{\prime})}$
$=\sum_{u}\sum_{y}y\frac{P(u,t^{\prime},y,x^{\prime})}{P(t^{\prime}|u)P(x^{\prime})}$
 [^10] The equation $P(x^{\prime}|u,t^{\prime},y)=P(x^{\prime}|u)$ is derived by from the conditional independence.
The second equality holds by Theorem 4, while the other equalities is grounded in basic derivations and property of probability.
On the other hand, as the expectation function of random variables $T,X,Y,U$,
$E[\frac{Y\mathbf{1}_{T=t^{\prime},X=x^{\prime}}}{P(T=t^{\prime}|U)P(X=x^{\prime})}]$ $=\sum_{u,t,y,x}\frac{y\mathbf{1}_{t=t^{\prime},x=x^{\prime}}}{P(t^{\prime}|u)P(x^{\prime})}P(u,t,y,x)$
$=\sum_{u,y}\sum_{t}\sum_{x}\frac{y\mathbf{1}_{t=t^{\prime}}\cdot\mathbf{1}_{x=x^{\prime}}}{P(t^{\prime}|u)P(x^{\prime})}P(u,t,y,x)$
$=\sum_{u,y}\sum_{t}\frac{y\mathbf{1}_{t=t^{\prime}}}{P(t^{\prime}|u)P(x^{\prime})}P(u,t,y,x^{\prime})$
$=\sum_{u}\sum_{y}\frac{y}{P(t^{\prime}|u)P(x^{\prime})}P(u,t^{\prime},y,x^{\prime})$
Combining the above two formulas, we complete our proof. ∎
Theorem 8.
For a DiscoSCM with discrete treatment $T$, feature $X$ and outcome $Y$:
$E[Y(t)|X=x,T=t,Y=y]=E[\frac{Y\mathbf{1}_{T=t,X=x}}{P(t,x)}\cdot\frac{P(y|t;U)}{P(y|t,x)}]$ (30)
where $x,t,y$ are the observed values of features, treatment and outcome.
Proof.
On one hand, we have:
$E[Y(t)|X=x,T=t,Y=y]$
$=E[E[Y(t)|X=x,T=t,Y=y,U]|X=x,T=t,Y=y]$
$=\sum_{u}\sum_{y^{*}}{y^{*}}P(y^{*}_{t}|x,t,y;u)P(u|x,t,y)$
$=\sum_{u}\sum_{y^{*}}{y^{*}}P({y^{*}}|{t};u)P(u|x,t,y)$
$=\sum_{u}\sum_{y^{*}}{y^{*}}\frac{P(y^{*}|t,u)P(t|u)P(u)}{P(t|u)}\cdot\frac{P(x,t,y|u)}{P(x,t,y)}$
$=\sum_{u}\sum_{y^{*}}y^{*}\frac{P(u,t,y^{*})}{P(t|u)}\cdot\frac{P(x|u,t,y^{*})}{P(x|u,t,y)}\cdot\frac{P(x,t,y|u)}{P(x,t,y)}$
$=\sum_{u}\sum_{y^{*}}y^{*}\frac{P(u,t,y^{*},x)}{P(t|u)P(x,t,y)}P(t,y|u)$
The second equality holds by Theorem 4, while the other equalities is grounded in basic derivations and property of probability. On the other hand, as the expectation function of random variables $T,X,Y,U$,
$E[\frac{Y\mathbf{1}_{T=t,X=x}}{P(t,x)}\cdot\frac{P(y|t;U)}{P(y|t,x)}]$ $=\sum_{u,t^{*},y^{*},x^{*}}\frac{y^{*}\mathbf{1}_{t=t^{*},x=x^{*}}}{P(t,x)}\cdot\frac{P(y|t;u)}{P(y|t,x)}P(u,t^{*},y^{*},x^{*})$
$=\sum_{u,y^{*}}\sum_{t^{*}}\sum_{x^{*}}\frac{y^{*}\mathbf{1}_{t=t^{*}}\cdot\mathbf{1}_{x=x^{*}}}{P(t,x)}P(u,t^{*},y^{*},x^{*})\cdot\frac{P(y|t;u)}{P(y|t,x)}$
$=\sum_{u,y^{*}}\sum_{t^{*}}\frac{y^{*}\mathbf{1}_{t=t^{*}}}{P(t,x)}P(u,t^{*},y^{*},x)\cdot\frac{P(y|t;u)}{P(y|t,x)}$
$=\sum_{u}\sum_{y^{*}}\frac{y^{*}}{P(t,x)}P(u,t,y^{*},x)\cdot\frac{P(y|t;u)}{P(y|t,x)}$
Combining the above two formulas, we complete our proof. ∎